Victoria Enterprises expects earnings before interest and taxes (EBIT) next year of $2.4 million. Its depreciation and capitaNeed help answering this question. Thanks for the help!
Answer #1
FCFF = EBIT*(1-tax rate)+Depreciation- Capex- Chng. In Work. Cap.
FCFF = 2400000*(1-0.35)+289000-289000-47000
FCFF = 1513000
firm or enterprise value = FCF in 1 year/(WACC - growth rate)
Firm/enterprise value = 1513000/ (0.08 - 0.06)
Firm/enterprise value = 75650000